# StarStream on Gaia: Stream discovery and mass loss rate of globular clusters

### v1

@@ 5 Summary and discussion
Our validation tests verified the near-unity detection ratio for the high-quality sample, enabling us to obtain first-ever unbiased estimate of the stream density. Based on this density, we calculate the mass loss rate of their progenitor GCs (Fig. 5). However, the detection ratio has a log-normal scatter varying between 0.15 dex and 0.5 dex depending on the background density (Appendix A and Fig. 6), dominating the uncertainty of $\dot{M}$. As discussed in C25, next generation wide-field surveys are likely to reduce this scatter by providing either additional independent observables (such as metallicity and line-of-sight velocity) or lower photometry and proper motion uncertainties.
By fitting $\dot{M}$ as a multivariate power-law function of the GC mass $M_{\rm GC}$ and half-mass radius $r_{\rm h}$, we observe slightly positive slopes for both quantities with a large scatter $\approx 0.3$ dex. We note that this positive correlation is likely dominated by several “fluffy” GCs with low mass and large radius. These GCs show high $|\dot{M}|\approx 10\ {\rm M_{\odot}\,Myr^{-1}}$. They are consistent with the BH-rich scenario of M. Gieles et al. (2021), where the BH population enhances both the half-mass radius and mass loss rate near the end of the GC’s lifetime.
@@ Acknowledgments
We thank Monica Valluri, Eric Bell, and Colin Holm-Hansen for insightful discussions. We thank Davide Massari for sharing updated extinction measurements for select GCs. YC and OYG were supported in part by National Aeronautics and Space Administration through contract NAS5-26555 for Space Telescope Science Institute programs HST-AR-16614 and JWST-GO-03433. This research benefited from the Gravity in the Local Group conference hosted by the McWilliam’s Center for Cosmology and Astrophysics, Carnegie Mellon University.
Appendix A Proof of the unbiased estimate of mass loss rate
@@
$N_{\rm tracer}=\lambda\,f_{\rm detect}^{-1}N_{\rm detect}.$ (A1)
@@
Appendix B Low-quality sample
We provide the GC properties and number of detections $N_{\rm detect}$ for the low-quality sample in Table 2.

### v2

@@ V Summary and discussion
Our validation tests verified the near-unity detection ratio for the high-quality sample, enabling us to obtain first-ever unbiased estimate of the stream density. Based on this density, we calculate the mass loss rate of their progenitor GCs (Fig. 5). However, the detection ratio has a log-normal scatter varying between 0.15 dex and 0.5 dex depending on the background density (Appendix B and Fig. 6), dominating the uncertainty of $\dot{M}$. As discussed in C25, next generation wide-field surveys are likely to reduce this scatter by providing either additional independent observables (such as metallicity and line-of-sight velocity) or lower photometry and proper motion uncertainties.
By fitting $\dot{M}$ as a multivariate power-law function of the GC mass $M_{\rm GC}$ and half-mass radius $r_{\rm h}$, we observe the slopes for both quantities are consistent with zero with large scatters. We also notice several “fluffy” GCs with low mass and large radius. These GCs show high $|\dot{M}|\approx 10\ {\rm M_{\odot}\,Myr^{-1}}$. They are consistent with the BH-rich scenario of M. Gieles et al. (2021), where the BH population enhances both the half-mass radius and mass loss rate near the end of the GC’s lifetime.
@@ Acknowledgments
We thank Monica Valluri, Eric Bell, and Colin Holm-Hansen for insightful discussions. We thank Davide Massari for sharing updated extinction measurements for select GCs. We also thank the reviewer for suggestions and comments, which have improved the quality of this work. YC and OYG were supported in part by National Aeronautics and Space Administration through contract NAS5-26555 for Space Telescope Science Institute programs HST-AR-16614 and JWST-GO-03433. This research benefited from the Gravity in the Local Group conference hosted by the McWilliam’s Center for Cosmology and Astrophysics, Carnegie Mellon University.
Appendix A Special treatment for NGC 5024 and 5053
The method outlined in §II.1 assumes that there is no significant overlap between multiple streams. This assumption is not valid for the GC pair NGC 5024 and 5053, which reside near the Galactic pole and are separated by only $\sim 1^{\circ}$. Z. Yuan et al. (2020) suggested that the streams associated with these two clusters overlap substantially, although their radial velocities differ by $\sim 100\ {\rm km\,s^{-1}}$. To account for this complexity, we adopt a three-component mixture model to describe the probability density function with two distinct streams:
$p({\bm{x}})=f_{\rm s1}\,p_{\rm s1}({\bm{x}})+f_{\rm s2}\,p_{\rm s2}({\bm{x}})+(1-f_{\rm s1}-f_{\rm s2})\,p_{\rm bg}({\bm{x}}),$
where indices 1 and 2 correspond to the streams of NGC 5024 and 5053, respectively. The resulting model therefore introduces two free parameters, $f_{\rm s1}$ and $f_{\rm s2}$, which we fit jointly by maximizing the corresponding likelihood function.
Once we obtain the best-fit values of $f_{\rm s1}$ and $f_{\rm s2}$, we compute the probability that star $i$ belongs to stream 1 as
$P_{{\rm s1},i}\equiv\frac{f_{\rm s1}\,p_{\rm s1}({\bm{x}}_{i})}{f_{\rm s1}\,p_{\rm s1}({\bm{x}}_{i})+f_{\rm s2}\,p_{\rm s2}({\bm{x}}_{i})+(1-f_{\rm s1}-f_{\rm s2})\,p_{\rm bg}({\bm{x}}_{i})}.$
The probability $P_{{\rm s2},i}$ is computed in a similar manner. We classify stars with $P_{{\rm s1},i}+P_{{\rm s2},i}>P_{\rm th}$ as stream members. Among these stars, those with $P_{{\rm s1},i}>P_{{\rm s2},i}$ are assigned to stream 1, while the remainder are assigned to stream 2.
As described in §III.1, our search is restricted to stars within a $10^{\circ}$ radius of each GC. For the GC pair here, the simulated stream of NGC 5053 is shorter than $10^{\circ}$ and is therefore fully enclosed within the search radius centered on NGC 5024. Consequently, we search for both streams within $10^{\circ}$ around NGC 5053. In addition, because either GC passes through the stream of each other, we exclude stars within the tidal radii of both clusters when performing the joint fit.
We find that the likelihood function peaks at $f_{\rm s1}=2.29\times 10^{-4}$ (NGC 5024) and $f_{\rm s2}=4.57\times 10^{-4}$ (NGC 5053), with no significant degeneracy between the two parameters. This indicates that the presence of both streams is necessary to describe the overall stellar distribution in the surrounding region. This appears to be in tension with Z. Yuan et al. (2020), who classified several stars as belonging to a single, wide stream associated with both GCs. However, their stars are predominantly far from the two clusters ($10^{\circ}$), where the individual streams may already be dynamically mixed.
Appendix B Proof of the unbiased estimate of mass loss rate
@@
$N_{\rm tracer}=\lambda\,f_{\rm detect}^{-1}N_{\rm detect}.$ (B1)
@@
Appendix C Low-quality sample
We provide the GC properties and number of detections $N_{\rm detect}$ for the low-quality sample in Table 2. Note that NGC 6715, Arp 2, Ter 7, and Ter 8 are members of the Sgr stream (M. Bellazzini et al., 2020; E. Vasiliev et al., 2021, marked with asterisk symbols in the table). We may overestimate $N_{\rm detect}$ by selecting stars from the Sgr stream, which shares similar spatial, proper motion, color, and magnitude distributions.